OpenStudy (anonymous):

Need help with LaPlace transform y''-y=5*sin2t when given y(0)=0, y'(0)=6

OpenStudy (anonymous):

I've gotten to \[2^2L(y)-6-y=5*\sin2t\]

OpenStudy (anonymous):

but I think I'm stuck... that should say s^2 not 2^2

OpenStudy (anonymous):

\[s^2L(y)-6-y=5*\sin(2t)\]

OpenStudy (anonymous):

do you have the value of y'(0)?

OpenStudy (anonymous):

yep y'(0)=6

OpenStudy (anonymous):

I messed that all up.... it should be\[s^2L(y)-6-L(y)=10/s^2+4\]

OpenStudy (anonymous):

yeah this is right

OpenStudy (anonymous):

Ok so what do I do from there?

OpenStudy (anonymous):

try to solve for L(y)

OpenStudy (anonymous):

I thikn I would go \[s^2(L(y)-1)-6=10/s^2+4\]

OpenStudy (anonymous):

err

OpenStudy (anonymous):

ok I prefer to write L(y) as Y(s)

OpenStudy (anonymous):

\[L(y)(s^2-1)-6=10/s^2+4\]

OpenStudy (anonymous):

it should be : \[Y(s)(s^2-1)={10 \over s^2+4} +6\]

OpenStudy (anonymous):

then you divide each side by s^2-1 and then look up the transform on a table?

OpenStudy (anonymous):

=>6s^2+34 = Y{s^2 - 1}

OpenStudy (anonymous):

Im confused on where you got that thinker

OpenStudy (anonymous):

take the LCM :)

OpenStudy (anonymous):

= \[Y(s)={6s^2+34 \over s^2+4} . {1 \over s^2 -1}\]

OpenStudy (anonymous):

sry :) missed the denominator :)

OpenStudy (anonymous):

are you follwoing Scotty?

OpenStudy (anonymous):

now i guess u may use partial fractions to solve further

OpenStudy (anonymous):

1 sec I think so im writing it out

OpenStudy (anonymous):

take your time.. I will be right back

OpenStudy (anonymous):

try a way to write it back in time domain

OpenStudy (anonymous):

Yep I got it, you were adding 6 into there.

OpenStudy (anonymous):

Can I look up the laplace transform inverse for each and then combine them?

OpenStudy (anonymous):

i guess its possible when there's addition there..here we have multiplication of terms...can u use partial fractions?

OpenStudy (anonymous):

Forgot how to do those, you're talking about breaking it up right?

OpenStudy (anonymous):

take , say, (A/s^ + 4)+(b/s^2 -1)

OpenStudy (anonymous):

sry it is (A/s^2 + 4)+(b/s^2 -1)

OpenStudy (anonymous):

mm ok I got it now. Thanks guys... I'm sure I'm gonna post another one here in a min lol

OpenStudy (anonymous):

I am back

OpenStudy (anonymous):

I think there is something wrong

OpenStudy (anonymous):

no I just ran what we had through my solver and it came up with the correct answer...

OpenStudy (anonymous):

I mean the partial fractions

OpenStudy (anonymous):

whats that?? u know it was 2 years back i studies Laplace transforms :):)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Oh well the laplace transform before you get to the fractions are correct :)

OpenStudy (anonymous):

^^ yeah I know :)

OpenStudy (anonymous):

omg!! did i forget solving these?? lol

OpenStudy (anonymous):

if that's all you want, then it's ok.. but if you want to fin the final solution for the differential equation by yourself, then you need to do the partial fractions

OpenStudy (anonymous):

right

OpenStudy (anonymous):

the point is that you want to write the expression we have for Y(s) in a form that is easy to be transformed back to time domain

OpenStudy (anonymous):

I didn't learn the notation in form of Y(s) which seems to be the way most people learn it. Very annoying.

OpenStudy (anonymous):

was my suggestion of no help ??

OpenStudy (anonymous):

No you helped thinker. Thank you :)

OpenStudy (anonymous):

it doesn't matter what notation you use as long as you understand it.

OpenStudy (anonymous):

thnx :) @scott